OpenStudy (kobeni-chan):

Need help with negative rational exponents - will give medal! 1/6-3 = 1/1/63 = 1 ÷ 1/63 = 1 .(multiplied by) 63/1 = 216 <--This is the example I was given. I understand all of these steps except the last one... to me, 63/1 is 63, then multiplied by 1 is still 63...so how did it turn out to be 216?

OpenStudy (tkhunny):

You MUST get better notation. Very VERY difficult to understand what you are doing. Use the carat to indicate exponentiation. "six to the 3rd" can be 6^3 "six to the -3rd" can be 6^(-3) \(\dfrac{1}{6^{-3}} = \dfrac{1}{\dfrac{1}{6^{3}}}\) Okay, first, I would suggest you NEVER do this. Save yourself a TON of trouble and possibility for error like this... \(\dfrac{1}{6^{-3}} = 6^{3}\) - - Done! 6^3 = 6*6*6 = 36*6 = 216 Another example: \(\dfrac{5^{-2}}{4^{-5}} = \dfrac{4^{5}}{5^{2}}\) -- Absolutely no need to mess around with those wonderfully- and unnecessarily-complicated fractions.

OpenStudy (kobeni-chan):

oh! that made it a lot simpler lol

OpenStudy (triciaal):

@tkhunny I am so glad someone could translate!!!